OpenStudy (anonymous):

An ordinary glass is filled to the brim with 322.6mL of water at 100.0 degrees C . If the temperature decreased to 20.0 C , how much water could be added to the glass? For the change in volume I multiplied coef. of volumetric thermal expansion (207*10^-6)(100-20 C)*(322.6mL). I got 5.34mL. I did convert celcius to Kelvin so the units cancel out. But either way you still get the same answer. The answer I got was wrong?? What do I need to do to fix this??

OpenStudy (anonymous):

I get 5.52mL using 214*10-6 as you thermal expansion coefficient.

OpenStudy (anonymous):

that is not the answer either :/

OpenStudy (anonymous):

how about 18,23 ml?

OpenStudy (anonymous):

and here is table, i think you need to use 100°C coefficient, if not try average between 20 and 100 or somethnig like that: http://physchem.kfunigraz.ac.at/sm/Service/Water/H2Othermexp.htm

OpenStudy (anonymous):

That is still is not the answer apparently I don't know what the right answer is supposed to be

OpenStudy (anonymous):

Volume of the glass = 322.6mL The Density variation with temperature is not leniar, so for large temperature differences it is better to refer to the density tables. Mass remains constant, M=Density*Volume D(20)*V(20)=D(100)*V(100)=Mass [V(20) == volume at 20deg celcius] V(20)=D(100)*V(100)/D(20) D(20)=998.2071 Kg/m3; D(100)=958.4 Kg/m3 (From Wikipedia) V(20)=958.4*322.6/998.2071 =309.73mL V extra = 322.6-309.73 =12.86mL

OpenStudy (anonymous):

@MartialArtGirl647 are u there?

OpenStudy (anonymous):